# Myers–Steenrod theorems for metric and singular Riemannian foliations

### v1

@@ 2.3. Submetries from proper metric spaces
Let $p\colon X\to Y$ be a submetry between two metric spaces. Two points $x_{1},x_{2}\in X$ are near (with respect to $p$) if $d(x_{1},x_{2})=d(p(x_{1}),p(x_{2}))$. The points $x_{1},x_{2}$ are near if and only if they realize the distance between the leaves $L_{x_{1}}$ and $L_{x_{2}}$ in $X$.
Consider a length space $X$, a metric space $Y$, and a submetry $p\colon X\to Y$. Let $\gamma$ be a geodesic in $Y$ through a point $y\in Y$. Given $x\in p^{-1}(y)$, there exists a geodesic $\tilde{\gamma}$ through $x$ that is mapped by $p$ isometrically onto $\gamma$. We call $\tilde{\gamma}$ the horizontal lift of $\gamma$. A geodesic in $X$ is horizontal if it is mapped by $p$ isometrically onto a geodesic in $Y$. A shortest path between two points $x_{1},x_{2}\in X$ is horizontal if and only if the points $x_{1},x_{2}$ are near (see [29, p. 21]).
Theorem 2.3 ([29, Theorem 7.2]).
@@
2.4. Submetries between Alexandrov spaces
@@
Let $(X,d)$ be a metric space. For each point $x\in X$, we let
$\operatorname{rad}_{x}=\sup_{y\in X}d(x,y),$
and define the radius of $X$, denoted by $\operatorname{rad}(X)$, by
$\mathrm{rad}(X)=\inf\{\mathrm{rad}_{x}\mid x\in X\}.$
Let $\Sigma$ be an $n$-dimensional Alexandrov space with curvature bounded below by $1$. Following [29, Definition 2.5], we will say that $\Sigma$ is Riemannian if it is isometric to the unit round $n$-sphere $S^{n}$. If $\operatorname{vol}(\Sigma)>(1-\rho)\operatorname{vol}(S^{n})$ for a fixed and sufficiently small positive real number $\rho=\rho(n)$, we refer to $\Sigma$ as extremely thick. If there exist $(n+2)$ points $v_{i}\in\Sigma$ with $d(v_{i},v_{j})>\pi/2$, we call $\Sigma$ thick. Additionally, we say $\Sigma$ is round if $\mathrm{rad}(\Sigma)>\pi/2$. By Grove and Petersen’s radius sphere theorem, every round Alexandrov space is homeomorphic to a sphere (see [23]). Given $A\subset\Sigma$, we let $\operatorname{Pol}(A)$ be the set of all $v\in\Sigma$ with $d(v,A)\geqslant\pi/2$, and refer to it as the polar set of $A$. Two points $v,w\in\Sigma$ are antipodal if $d(v,w)=\pi$. Toponogov’s comparison theorem implies that $\operatorname{Pol}(A)$ is a totally convex subset of $\Sigma$. Hence, $\operatorname{Pol}(A)$ is an Alexandrov space with $\operatorname{curv}\geqslant 1$.
Let $X$ and $Y$ be metric spaces and let $CX$ and $CY$ be their respective topological cones. A map $f\colon CX\to CY$ is homogeneous if
$f([t,x])=[t,f(x)]$
for all $t\in[0,\infty)$, and all $x\in X$. As mentioned in Section 2.2, if $\Sigma$ is an $n$-dimensional Alexandrov space with $\operatorname{curv}\geq 1$, then $C\Sigma$ may be endowed with a metric $d$ such that $(C\Sigma,d)$ is an $(n+1)$-dimensional Alexandrov space of non-negative curvature, the euclidean cone of $\Sigma$. We will refer to points in $C\Sigma$ as vectors. Given a vector $v=[t,u]$ in $(C\Sigma,d)$, we will refer to $t$ as the magnitude of $v$ and will set $|v|=t$. Let $p\colon C\Sigma\to C\mathrm{T}$ be a submetry between euclidean cones. A vector $h\in C\Sigma$ is horizontal (with respect to $p$) if $|p(h)|=|h|$.
Proposition 2.4 ([29, Proposition 6.4 and Lemma 6.5]).
Let $\Sigma$ and $\mathrm{T}$ be Alexandrov spaces with $\operatorname{curv}\geqslant 1$ and let $C\Sigma$, and $C\mathrm{T}$ be their respective euclidean cones. If $p\colon C\Sigma\to C\mathrm{T}$ is a homogeneous submetry, then the following assertions hold:
@@
Let $V$ and $H$ be as in Proposition 2.4. Given $v\in V$ with $|v|=1$, consider the set
$H^{v}=\{h\in X\mid d(h,v)=\pi/2\}\subset\operatorname{Pol}(V)=H.$
The set $\widetilde{H}=\cap_{v\in V}H^{v}$ is the set of all points $h\in H$ that have distance equal to $\pi/2$ to any point in $V$. Since $H$ and $V$ are totally convex subsets of $\Sigma$, they are Alexandrov spaces of $\operatorname{curv}\geqslant 1$. Thus their spherical join $H*V$ is again an Alexandrov space with $\operatorname{curv}\geq 1$. Moreover, there exists a $1$-Lipschitz map $P_{H}\colon\Sigma\to H\ast V$ for which the following assertions are equivalent:
Proposition 2.5 ([29, Proposition 6.14 ]).
@@
If the submetry $P_{H}\colon\Sigma\to H\ast V$ satisfies any of the conditions in Proposition 2.5, we say that the set $H\subset\Sigma$ is almost spanning. Moreover, if $P_{H}$ is an isometry, we say that $H$ is spanning. We call a homogeneous submetry $p\colon C\Sigma\to C\mathrm{T}$ (almost) spanning if the horizontal set $H\subset\Sigma$ is (almost) spanning.
By considering the euclidean cones over the spaces $\Sigma$, $H\ast V$, $H$, and $\mathrm{T}$, and the maps between them discussed above, we may represent a submetry $p\colon C\Sigma\to C\mathrm{T}$ as a composition
[…]
where $\mathrm{pr}_{CH}$ is the projection onto the first factor of the metric product $CH\times CV$, and the other two maps are submetries.
Lemma 2.6 ([29, Lemma 6.15]).
Let $p\colon C\Sigma\to C\mathrm{T}$ be a homogeneous submetry and let $H\subset\Sigma$ be the set of horizontal vectors. If the restriction ${\left.\kern-1.2ptp\right|_{H}}\colon H\to\mathrm{T}$ is an isometry, then $H$ is almost-spanning.
@@
Proposition 2.7 ([29, Lemma 6.16]).
@@
We now recall the definition of differentiability for Lipschitz functions between Alexandrov spaces (cf. [29, Section 3]). Given a metric space $(X,d)$ and a real number $r>0$, we denote by $rX$ the metric space $(X,rd)$. Let $\{(X_{i},x_{i})\}_{i=1}^{\infty}$ and $(X,x)$ be pointed proper metric spaces. The sequence $\{(X_{i},x_{i})\}$ converges in the pointed Gromov–Hausdorff sense to $(X,x)$ if, for each $R>0$, the sequence of closed balls $\{B_{R}(x_{i})\}$ converges to $B_{R}(x)$ in the Gromov–Hausdorff sense. The following theorem gives an alternative characterization of the tangent cone of an Alexandrov space to the one in Section 2.2.
Theorem 2.8 ([9, Theorem 7.8.1]).
Let $X$ be an Alexandrov space, fix $x\in X$, and let $T_{x}X=C\Sigma_{x}(X)$ be the tangent cone to $X$ at $x$. Then the pointed metric spaces $(\lambda X,x)$ converge in the pointed Gromov-Hausdorff sense to $(T_{x}X,0)$ as $\lambda\to\infty$.
A map $f\colon(X,d_{X})\to(Y,d_{Y})$ between metric spaces is Lipschitz if there exists a real number $K>0$ such that $d_{Y}(f(x_{1}),f(x_{2}))\leq Kd_{X}(x_{1},x_{2})$ for any $x_{1},x_{2}\in X$. Let $X$, and $Y$ be Alexandrov spaces, $U$ an open subset of $X$, and fix $x\in U$. Let $f\colon U\to Y$ be a Lipschitz map and let $\{r_{j}\}$ be a sequence of positive real numbers such that $r_{j}\to 0$ as $j\to\infty$. Then there exists a limit map
@@
We say that $f$ is differentiable at $x$, if the map $f_{(r_{j})}$ is independent of the sequence $\{r_{j}\}$. We call this uniquely defined Lipschitz function the differential of $f$ at $x$, and denote it by $df_{x}\colon T_{x}X\to T_{f(x)}Y$. If the differential of $f$ exists at every point $x\in U$, then we say that $f$ is differentiable. If $f$ is differentiable at $x\in U$, then its differential $df_{x}$ is homogeneous. That is, $df_{x}(tv)=tdf(v)$ for each real number $t\geq 0$ and each $v\in T_{x}X$.
@@
Moreover, a submetry $p\colon X\to Y$ from an Alexandrov space has a well defined differential.
Proposition 2.10 ([29, Proposition 5.1]).
Let $p\colon X\to Y$ be a submetry from an Alexandrov space $X$ to $Y$ a metric space. Then $p$ is differentiable and each differential $df_{x}\colon T_{x}X\to T_{p(x)}Y$ is a homogeneous submetry.
Let $p\colon X\to Y$ be a submetry from an Alexandrov space $X$ to $Y$ a metric space. A vector $v\in T_{x}X$ is vertical if $df(v)=0$. We call $v$ horizontal if $|Df(v)|=|v|$. From the homogeneity of the differential, one may show that the set of vertical vectors forms a subcone $CV_{x}$ of $T_{x}X$, where $V_{x}\subset\Sigma_{x}X\subset T_{x}X$. Similarly, the set of horizontal vectors forms a subcone $CH_{x}$ of $T_{x}X$, where $H_{x}\subset\Sigma_{x}X\subset T_{x}X$.
A submetry $p\colon X\to Y$ between arbitrary Alexandrov spaces is regular at a point $x\in X$ if the homogeneous submetry $dp_{x}\colon T_{x}X\to T_{p(x)}Y$ is regular. We call the leaf $L_{x}$ a regular leaf. Similarly, $p\colon X\to Y$ is (almost) spanning at $x\in X$ if $dp_{x}\colon T_{x}X\to T_{p(x)}Y$ is (almost) spanning. A point $x\in X$ is round if $\Sigma_{x}(X)$ is a round space. The following corollary follows immediately from Proposition 2.7.
@@
Theorem 2.12 ([8, Theorem 10.8.5] and [8, Corollary 10.9.13]).
If $X$ is an $n$-dimensional Alexandrov space, then the set of points whose space of directions is a unit round sphere (or, equivalently, whose tangent cone is Euclidean space) is dense in $X$.
Recall that, given a submetry $p\colon X\to Y$ between metric spaces, the intrinsic metric on the fiber $L=p^{-1}(y)$ is locally Lipschitz-equivalent to the induced metric on $L$ (see Theorem 2.3). Thus, when $X$ is an Alexandrov space and $p$ is a proper submetry, any fiber $L$ is an Alexandrov space.
The following proposition allows us to identify, for any $x\in X$, the space of vertical directions $V_{x}\subset\Sigma_{x}X$ with $\Sigma_{x}(L)$, the space of directions of the leaf $L$ trough $x$.
Proposition 2.13 ([29, Proposition 5.2]).
Let $p\colon X\to Y$ be a proper submetry between Alexandrov spaces. Then, for any fiber $L=p^{-1}(y)$ and any point $x\in L$, the tangent cone $T_{x}L$ is the cone $CV_{x}$.
@@
Lemma 2.14 (Existence of geodesic lifts: [25, Lemma 1] and [4, Lemma 2.1]).
Let $f\colon X\to Y$ be a submetry between Alexandrov spaces. For any geodesic $\alpha\colon[0,a]\to Y$, and any $x\in f^{-1}(\alpha(0))$, there is a geodesic lift $\tilde{\alpha}\colon[0,a]\to X$ of $\alpha$ starting at $x$, with length equal to that of $\alpha$. Moreover, if $\alpha$ is the geodesic $\alpha$ between $\alpha(0)$ and $\alpha(a)$ is unique, then the lift $\tilde{\alpha}$ is unique.
2.5. Holonomy maps of a submetry between Alexandrov spaces
Let $f\colon X\to Y$ be a submetry between Alexandrov spaces. Assume we have a geodesic $\alpha\colon[0,b]\to Y$ with $\alpha(0)=y$, $\alpha(b)=z$, and initial direction $w\in\Sigma_{y}Y$. By [29, Lemma 5.4], for $x\in f^{-1}(y)$ fixed, for each $h\in H_{x}$ with $df_{x}(h)=w$, there exists a unique geodesic $\tilde{\alpha}_{h}\colon[0,b]\to X$ starting at $x$ with initial direction $h$, such that $f\circ\tilde{\alpha}_{h}=\alpha$. If there exists a unique direction $\tilde{w}\in H_{x}$ such that $df_{x}(\tilde{w})=w$, then we have a well-defined map $p_{\alpha}\colon f^{-1}(y)\to f^{-1}(z)$ given by
$p_{\alpha}(x)=\alpha_{\tilde{w}}(b).$
This map is called the holonomy map of $\alpha$.
Consider a geodesic $\alpha\colon[0,b]\to Y$ such that, for each $x\in f^{-1}(\alpha(0))$, the initial direction of $\alpha$ has a unique lift in $H_{x}$ and $p_{\alpha}$ is well defined. As observed in [29, Section 7.3], for $\tilde{z}\in f^{-1}(\alpha(b))$, the set $p^{-1}_{\alpha}(\tilde{z})$ corresponds to the endpoints of all geodesics of $X$ starting at $\tilde{z}$ which are horizontal lifts of the geodesic $\alpha^{-1}(t)=\alpha(b-t)$ and whose initial direction is a horizontal lift of the initial direction of $\alpha^{-1}$.

### v3

@@ 3. $(3)$
Let $p\colon X\to Y$ be a submetry between two metric spaces. Two points $x_{1},x_{2}\in X$ are near ((with respect to $p)$ if $d(x_{1},x_{2})=d(p(x_{1}),p(x_{2}))$. The points $x_{1}$, $x_{2}$ are near if and only if they realize the distance between the leaves $L_{x_{1}}$ and $L_{x_{2}}$ in $X$.
Consider a length space $X$, a metric space $Y$, and a submetry $p\colon X\to Y$. Let $\gamma$ be a geodesic in $Y$ through a point $y\in Y$. Given $x\in p^{-1}(y)$, there exists a geodesic $\tilde{\gamma}$ through $x$ that is mapped by $p$ isometrically onto $\gamma$. We call $\tilde{\gamma}$ the horizontal lift of $\gamma$. A geodesic in $X$ is horizontal if it is mapped by $p$ isometrically onto a geodesic in $Y$. A shortest path between two points $x_{1},x_{2}\in X$ is horizontal if and only if the points $x_{1}$, $x_{2}$ are near (see [36, p. 21]).
Theorem 2.3 ([36, Theorem 7.2]).
@@ 2. $(2)$
2.4 Submetries between Alexandrov spaces
@@
Recall that a map $f\colon(X,d_{X})\to(Y,d_{Y})$} between metric spaces is Lipschitz if there exists a real number K>0{K>0} such that $d_{Y}(f(x_{1}),f(x_{2}))\leq Kd_{X}(x_{1},x_{2})$ for any $x_{1},x_{2}\in X$. In this case, we say $f$ is $K$-Lipschitz.
Let $\Sigma$ be an $n$-dimensional Alexandrov space with curvature bounded below by $1$. Given $A\subset\Sigma$, we let
$\operatorname{Pol}(A)=\{v\in\Sigma\mid d_{\Sigma}(v,A)\geq\pi/2\}$
and refer to it as the polar set of $A$. Two points $v,w\in\Sigma$ are antipodal if $d_{\Sigma}(v,w)=\pi$. Toponogov’s comparison theorem implies that $\operatorname{Pol}(A)$ is a totally convex subset of $\Sigma$. Hence, $\operatorname{Pol}(A)$ is an Alexandrov space with $\operatorname{curv}\geq 1$.
Let $X$ and $Y$ be metric spaces, and let $CX$ and $CY$ be their respective topological cones. A map $f\colon CX\to CY$ is homogeneous if $f([t,x])=[t,f(x)]$ for all $t\in[0,\infty)$, and all $x\in X$. As mentioned in Section 2.2, if $\Sigma$ is an $n$-dimensional Alexandrov space with $\operatorname{curv}\geq 1$, then $C\Sigma$ may be endowed with a metric $d$ such that $(C\Sigma,d)$ is an $(n+1)$-dimensional Alexandrov space of non-negative curvature. We call $(C\Sigma,d)$ the Euclidean cone over $\Sigma$. We will refer to points in $C\Sigma$ as vectorsand to points in $\Sigma$ as directions. Given a vector $v=[t,u]$ in $(C\Sigma,d)$, we will refer to $t$ as the magnitude of $v$ and will set $|v|=t$. Let $p\colon C\Sigma\to C\mathrm{T}$ be a submetry between Euclidean cones. A vector $h\in C\Sigma$ is horizontal ((with respect to $p)$ if $|p(h)|=|h|$. A subset $A$ of an Alexandrov space $X$ is totally convex if, for any two points $x,y\in A$, every (minimizing) geodesic joining $x$ and $y$ is contained in $A$. If $\operatorname{curv}(X)\geq 1$, we require this only for pairs with $d(x,y)<\pi$. In particular, $S^{0}$ is always a totally convex subset in any round unit sphere.
Proposition 2.4 ([36, Proposition 6.4 and Lemma 6.5]).
Let $\Sigma$ and $\mathrm{T}$ be Alexandrov spaces with $\operatorname{curv}\geq 1$ and let $C\Sigma$, and $C\mathrm{T}$ be their respective Euclidean cones. If $p\colon C\Sigma\to C\mathrm{T}$ is a homogeneous submetry, then the following assertions hold:
@@ 3. $(3)$
Let $V$ and $H$ be as in Proposition 2.4. Given $v\in V$, consider the set
$H^{v}=\{h\in\Sigma\mid d_{\Sigma}(h,v)=\pi/2\}.$
Since $v\in V\subset\Sigma$ and $d_{\Sigma}(h,v)=\pi/2$ implies $d_{\Sigma}(h,V)\geq\pi/2$, we have $H^{v}\subset\operatorname{Pol}(V)=H$. Thus, every $h\in H^{v}$ is a horizontal direction.
The set $\widetilde{H}=\bigcap_{v\in V}H^{v}$ is the set of all points $h\in H$ that have distance equal to $\pi/2$ to any point in $V$. Since $H$ and $V$ are totally convex subsets of $\Sigma$, they are Alexandrov spaces of $\operatorname{curv}\geq 1$}. Thus, their spherical join HVH*V is again an Alexandrov space with $\operatorname{curv}\geq 1$.
Moreover, there exists a $1$-Lipschitz map $P_{H}\colon\Sigma\to H\ast V$ for which the following assertions are equivalent.
Proposition 2.5 ([36, Proposition 6.14 ]).
@@
If the $1$-Lipschitz map $P_{H}\colon\Sigma\to H\ast V$ satisfies any of the conditions in Proposition 2.5, we say that the set $H\subset\Sigma$ is almost spanning. Moreover, if $P_{H}$ is an isometry, we say that $H$ is spanning. We call a homogeneous submetry $p\colon C\Sigma\to C\mathrm{T}$ ((almost)) spanning if the horizontal set $H\subset\Sigma$ is (almost) spanning.
By considering the Euclidean cones over the spaces $\Sigma$, $\ast V, HH$, and $\mathrm{T}$, and the maps between them discussed above, we may represent a homogeneous submetry $p\colon C\Sigma\to C\mathrm{T}$ as a composition
$C\Sigma\overset{CP_{H}}{\longrightarrow}CH\times CV\overset{\mathrm{pr}_{CH}}{\longrightarrow}CH\overset{p}{\longrightarrow}C\mathrm{T},$
where $\mathrm{pr}_{CH}$ is the projection onto the first factor of the metric product $CH\times CV$, and observe that $\mathrm{pr}_{CH}$ and $p$ are submetries.
Lemma 2.6 ([36, Lemma 6.15]).
Let $p\colon C\Sigma\to C\mathrm{T}$ be a homogeneous submetry and let $H\subset\Sigma$ be the set of horizontal directions. If the restriction ${\left.\kern-1.2ptp\right|_{H}}\colon H\to\mathrm{T}$ is an isometry, then $H$ is almost-spanning.
@@
Following [36, Definition 2.5], we will say that $\Sigma$ is Riemannian if it is isometric to the unit round $n$-sphere $S^{n}$. If $\operatorname{vol}(\Sigma)>(1-\rho)\operatorname{vol}(S^{n})$ for a fixed and sufficiently small positive real number $\rho=\rho(n)$, we refer to $\Sigma$ as extremely thick. If there exist $n+2$ points $v_{i}\in\Sigma$} with $d(v_{i},v_{j})>\pi/2$}, we call $\Sigma$ thick. Additionally, we say $\Sigma$ is round if $\mathrm{rad}(\Sigma)>\pi/2$, where $\operatorname{rad}(\Sigma)$, the radius of $\Sigma$, is given by
$\mathrm{rad}(\Sigma)=\inf_{v\in\Sigma}\sup_{w\in\Sigma}d(v,w).$
By Grove and Petersen’s radius sphere theorem, every round Alexandrov space is homeomorphic to a sphere (see [28]).
Proposition 2.7 ([36, Proposition 6.16]).
@@ 2. $(2)$
We now recall the definition of differentiability for Lipschitz functions between Alexandrov spaces (cf. [36, Section 3]). Given a metric space $(X,d)$ and a real number $r>0$, we denote by $rX$ the metric space $(X,rd)$. Let $\{(X_{i},x_{i})\}_{i=1}^{\infty}$ and $(X,x)$ be pointed proper metric spaces. The sequence $\{(X_{i},x_{i})\}$ converges in the pointed Gromov–Hausdorff sense to $(X,x)$ if, for each R>0{R>0}, the sequence of closed balls $\{B_{R}(x_{i})\}$ converges to $B_{R}(x)$ in the Gromov–Hausdorff sense. The following theorem gives an alternative characterization of the tangent cone of an Alexandrov space to the one in Section 2.2.
Theorem 2.8 ([12, Theorem 7.8.1]).
Let $X$ be an Alexandrov space, fix $x\in X$, and let $T_{x}X=C\Sigma_{x}(X)$ be the tangent cone to $X$ at $x$. Then the pointed metric spaces $(\lambda X,x)$ converge in the pointed Gromov–Hausdorff sense to $(T_{x}X,0)$ as $\lambda\to\infty$.
Let $X$, and $Y$ be Alexandrov spaces, $U$ an open subset of $X$, and fix $x\in U$. Let $f\colon U\to Y$ be a Lipschitz map and let $\{r_{j}\}$ be a sequence of positive real numbers such that $r_{j}\to 0$ as $j\to\infty$}. Then there exists a limit map
@@
We say that $f$ is differentiable at $x$, if the map $f_{(r_{j})}$ is independent of the sequence $\{r_{j}\}$. We call this uniquely defined Lipschitz function the differential of $f$ at $x$, and denote it by ${\rm d}f_{x}\colon T_{x}X\to T_{f(x)}Y$. If the differential of $f$ exists at every point $x\in U$, then we say that $f$ is differentiable. If $f$ is differentiable at $x\in U$, then its differential ${\rm d}f_{x}$ is homogeneous. That is, ${\rm d}f_{x}(tv)=t\,{\rm d}f(v)$ for each real number $t\geq 0$ and each $v\in T_{x}X$.
@@
Moreover, a submetry $p\colon X\to Y$ from an Alexandrov space has a well-defined differential.
Proposition 2.10 ([36, Proposition 5.1]).
Let $p\colon X\to Y$ be a submetry from an Alexandrov space $X$ to $Y$ a metric space. Then $p$ is differentiable and each differential ${\rm d}f_{x}\colon T_{x}X\to T_{p(x)}Y$ is a homogeneous submetry.
Let $p\colon X\to Y$ be a submetry from an Alexandrov space $X$ to $Y$ a metric space. A vector $v\in T_{x}X$ is vertical if ${\rm d}f_{x}(v)=0$. We call $v$ horizontal if $|{\rm d}f_{x}(v)|=|v|$. From the homogeneity of the differential, one may show that the set of vertical vectors forms a subcone $CV_{x}$ of $T_{x}X$, where $V_{x}\subset\Sigma_{x}X\subset T_{x}X$. Similarly, the set of horizontal vectors forms a subcone $CH_{x}$ of $T_{x}X$, where $H_{x}\subset\Sigma_{x}X\subset T_{x}X$.
A submetry $p\colon X\to Y$ between arbitrary Alexandrov spaces is regular at a point $x\in X$ if the homogeneous submetry ${\rm d}p_{x}\colon T_{x}X\to T_{p(x)}Y$ is regular. We call the leaf $L_{x}$ a regular leaf. Similarly, $p\colon X\to Y$ is ((almost)) spanning at $x\in X$ if ${\rm d}p_{x}\colon T_{x}X\to T_{p(x)}Y$ is (almost) spanning. A point $x\in X$ is round if $\Sigma_{x}(X)$ is a round space. The following corollary follows immediately from Proposition 2.7.
@@
Theorem 2.12 ([11, Theorem 10.8.5] and [11, Corollary 10.9.13]).
If $X$ is an $n$-dimensional Alexandrov space, then the set of points whose space of directions is a unit round sphere ((or, equivalently, whose tangent cone is Euclidean space)) is dense in $X$.
Recall that, given a submetry $p\colon X\to Y$ between metric spaces, the intrinsic metric on the fiber $L=p^{-1}(y)$ is locally Lipschitz-equivalent to the induced metric on $L$ (see Theorem 2.3).
The following proposition allows us to identify, for any $x\in X$, the space of vertical directions $V_{x}\subset\Sigma_{x}X$} with $\Sigma_{x}(L)$, the space of directions of the leaf $L$ through $x$.
Proposition 2.13 ([36, Proposition 5.2]).
Let $p\colon X\to Y$ be a submetry between Alexandrov spaces. Then, for any fiber $L=p^{-1}(y)$ and any point $x\in L$, the tangent cone $T_{x}L$ is the cone $CV_{x}$.
@@
Lemma 2.14 (existence of geodesic lifts: [30, Lemma 1] and [8, Lemma 2.1]).
Let $f\colon X\to Y$ be a submetry between Alexandrov spaces. For any geodesic $\alpha\colon[0,a]\to Y$, and any $x\in f^{-1}(\alpha(0))$, there is a geodesic lift $\tilde{\alpha}\colon[0,a]\to X$ of $\alpha$ starting at $x$, with length equal to that of $\alpha$. Moreover, if the geodesic $\alpha$ between $\alpha(0)$ and $\alpha(a)$ is unique, then the lift $\tilde{\alpha}$ is unique.
2.5 Holonomy maps of a submetry between Alexandrov spaces
Let $f\colon X\to Y$ be a submetry between Alexandrov spaces. Assume we have a geodesic $\alpha\colon[0,b]\to Y$ with $\alpha(0)=y$, $\alpha(b)=z$, and initial direction $w\in\Sigma_{y}Y$. By [36, Lemma 5.4] for $x\in f^{-1}(y)$ fixed, for each $h\in H_{x}$ with ${\rm d}f_{x}(h)=w$, there exists a unique geodesic $\tilde{\alpha}_{h}\colon[0,b]\to X$ starting at $x$ with initial direction $h$, such that $f\circ\tilde{\alpha}_{h}=\alpha$. If there exists a unique direction $\tilde{w}\in H_{x}$ such that $df_{x}(\tilde{w})=w$, then we have a well-defined map $p_{\alpha}\colon f^{-1}(y)\to f^{-1}(z)$ given by $p_{\alpha}(x)=\tilde{\alpha}_{\tilde{w}}(b)$. This map is called the holonomy map of $\alpha$.
[…]